covert to vertex form by completing the square.
y=-1/2x^2-7x+5

Answer by lwsshak3(11628)

covert to vertex form by completing the square.
y=-1/2x^2-7x+5
y=-1/2(x^2+14x+49)+24.5+5
y=-1/2(x+7)^2+29.5